# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v1

@@ B.3 Deriving the Closed-Form Solution for Ours
$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2},\quad\text{s.t.}\,(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (41)
This is a weighted least squares problem subject to an equality constraint. To solve it, we first formulate the Lagrangian function, where $\boldsymbol{\Lambda}$ is the Lagrange multiplier:
$\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})=\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2}+\boldsymbol{\Lambda}^{\top}\left((\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}\right).$ (42)
We compute the gradient of the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Delta}$ and set it to zero, yielding the following equation for $\boldsymbol{\Delta}$:
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}^{\top}+2\boldsymbol{\Delta}\mathbf{P}\mathbf{P}^{\top}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}^{\top}=\mathbf{0}.$ (43)
@@
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}+2\boldsymbol{\Delta}\mathbf{P}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}=\mathbf{0}.$ (44)
Therefore, we can obtain the closed-form solution for $\boldsymbol{\Delta}\mathbf{P}$ from this equation:
$\boldsymbol{\Delta}\mathbf{P}=(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}.$ (45)
Next, we differentiate the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Lambda}$ and set it to zero:
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Lambda}}=(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (46)
@@
$(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})\mathbf{M}\mathbf{C}_{2}=\mathbf{0}.$ (47)
@@
$\frac{1}{2}\boldsymbol{\Lambda}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}.$ (48)
@@
$(\boldsymbol{\Delta}\mathbf{P})_{\text{SPEED}}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{Q}\mathbf{M},$ (49)
@@ Appendix C Implementation Details
Table 7: The evaluation setup for multi-concept erasure. This celebrity dataset contains an erasure set with 100 celebrities and a retain set with another 100 celebrities. We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases.
@@ C.1 Experimental Setup Details
[…]
@@
Multi-concept erasure. We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task. Following the experiment setup from [35], we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 [1] can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) [20]. This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities. The full list for both sets is presented in Table 7. We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases. In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
Implicit concept erasure. We adopt the same setting in [19] to erase nudity $\to$ ‘ ’ as the erasure set and ‘ ’ as the retain set. In evaluation, we generate images using all 4,703 prompts in I2P and use NudeNet [6] to identify nude content with the threshold of 0.6.
@@
Table 8: Full quantitative comparison of the few-concept erasure in erasing instances from Table 1. The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
@@
Table 9: Full quantitative comparison of the few-concept erasure in erasing styles from Table 2. The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
Concept Snoopy Mickey Spongebob Pikachu Hello Kitty
CS FID CS FID CS FID CS FID CS FID
SD v1.4 28.51 - 26.62 - 27.30 - 27.44 - 27.77 -
Erase Snoopy
CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
SPM 23.72 116.26 26.62 31.21 27.21 31.96 27.41 19.82 27.80 30.95
SPM w/o FT 23.72 116.26 26.55 43.03 26.84 42.96 27.38 25.95 27.71 42.53
Ours 23.50 108.51 26.67 23.41 27.31 24.64 27.42 16.81 27.82 21.74
Erase Snoopy and Mickey
CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
SPM 23.18 122.17 22.71 117.30 26.92 38.35 27.35 27.13 27.76 39.61
SPM w/o FT 26.66 83.11 25.64 75.46 26.73 42.30 27.50 28.19 27.75 41.19
Ours 23.58 103.62 23.62 83.70 27.24 29.67 27.39 22.51 27.78 28.23
Erase Snoopy and Mickey and Spongebob
CS ↓ FID ↑ CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓
SPM 22.86 125.66 22.08 123.20 20.92 153.36 27.50 37.51 27.63 46.63
SPM w/o FT 27.75 68.89 26.24 63.30 24.82 77.38 27.44 30.00 27.74 38.89
Ours 23.69 103.33 23.93 86.55 21.39 109.28 27.47 21.40 27.76 26.22
Table 10: Quantitative comparison with SPM and SPM w/o FT (Facilitated Transport). The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
C.2 Erasure Configurations
Implementation of previous works. In our series of three concept erasure tasks, we mainly compare against four methods: ConAblhttps://github.com/nupurkmr9/concept-ablation [29], MACE§§§https://github.com/Shilin-LU/MACE [35], RECEhttps://github.com/CharlesGong12/RECE [19], and UCEhttps://github.com/rohitgandikota/unified-concept-editing [18], as they achieve SOTA performance across different concept erasure tasks. All the compared methods are implemented using their default configurations from the corresponding official repositories. One exception is that for MACE when erasing 50 celebrities, since it doesn’t provide an official configuration and the preserve weight varies with the number of target celebrities, we set it to $1.2\times 10^{5}$ to ensure a consistent balance between erasure and preservation.
Implementation of SPEED. In line with previous methods [29, 35, 19, 18], we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment [21]. In contrast, we only edit the value matrices in the CA layers, as suggested by [62]. This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images [58]. In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig. 5 and 6 (further ablation comparison is provided in Appx. D.3). The augmentation times $N_{A}$ in Eq. 8 is set to 10 and the augmentation ranks $r$ in Eq. 6 is set to 1 as ablated in Appx. D.3. Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the eigenvectors corresponding to the eigenvalues below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following [14]. Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq. 11 to ensure invertibility following [18].
@@ Appendix D Additional Experiments
D.1 Full Comparison on Few-Concept Erasure
@@
We present full quantitative comparisons of few-concept erasure, including both CS and FID, in Table 8 and Table 9. Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
D.2 More Baselines
In this section, we compare against more methods because of the page limit in our main paper. Since our method focuses on improving prior preservation and multi-concept erasure performance, we mainly compare it with similar methods, other methods like ESD [17], FMN [66], and SLD [51] are omitted, as they fail to achieve satisfactory prior preservation proved by previous comaprisons [37, 35, 62]. The remaining comparable method is SPM******https://github.com/Con6924/SPM [37], which is proposed to improve prior preservation and can scale to multi-concept erasure tasks. Notable, SPM not only fine-tunes the model weights using LoRA [25] but also intervenes in the image generation process through Facilitated Transport. Specifically, this module dynamically adjusts the LoRA scale based on the similarity between the sampling prompt and the target concept. In other words, if the prompt contains the target concept or is highly relevant, this scale is set to a large value, whereas if there is little to no relevance, it is set close to 0, functioning similarly to a text filter. We argue that such a comparison with SPM is not fair since we only focus on modifying the model parameters, and therefore, we compare both the original SPM and SPM without Facilitated Transport (SPM w/o FT) for a fair comparison. In the latter version, the LoRA scale is set to 1 by default.
The quantitative comparative results are shown in Table 10. It can be seen that our method consistently achieves the best prior preservation compared to both SPM and SPM w/o FT. Even equipped with Facilitated Transport, our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty). This superiority amplifies as the number of target concepts increases. For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. SPM’s 15.68 ↑). Once removing the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (highlighted by red in Table 10 and illustrated in Fig. 8). This indicates that the success of SPM in multi-concept erasure relies on the Facilitated Transport module, which dynamically allocates the LoRA scales by calculating the similarity between the sampling prompt and each target concept. For example, when erasing Snoopy + Mickey + SpongeBob, if the sampling prompt is “a photo of Snoopy”, SPM will allocate a larger scale to Snoopy’s LoRA according to the text similarity. On the contrary, if the sampling prompt is “a photo of Pikachu” with the non-target concept, all three LoRA scales will be assigned lower values, thereby preserving the prior knowledge. We argue that this strategy of dynamically tuning the LoRA scales based on the sampling prompt similarity is vulnerable to attacks and easily bypassed, especially in white-box attack scenarios, where an attacker can reconstruct the erased concepts by simply modifying the code with extremely low attack costs, e.g., open-source T2I models like Stable Diffusion [1, 2].

### v2

@@ B.3 Deriving the Closed-Form Solution for SPEED
$\mathop{\text{min}}\limits_{\bm{\Delta}}\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\bm{\Delta}\mathbf{P}\|^{2},\quad\text{s.t.}\,(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (41)
This is a weighted least squares problem subject to an equality constraint. To solve it, we first formulate the Lagrangian function, where $\bm{\Lambda}$ is the Lagrange multiplier:
$\mathcal{L}(\bm{\Delta},\bm{\Lambda})=\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\bm{\Delta}\mathbf{P}\|^{2}+\bm{\Lambda}^{\top}\left((\bm{\Delta}\mathbf{P})\mathbf{C}_{2}\right).$ (42)
We compute the gradient of the Lagrangian function in Eq. 42 with respect to $\bm{\Delta}$ and set it to zero, yielding the following equation for $\bm{\Delta}$:
$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Delta}}=2\left((\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}^{\top}+2\bm{\Delta}\mathbf{P}\mathbf{P}^{\top}+\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}^{\top}=\mathbf{0}.$ (43)
@@
$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Delta}}=2\left((\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}+2\bm{\Delta}\mathbf{P}+\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}=\mathbf{0}.$ (44)
Therefore, we can obtain the closed-form solution for $\bm{\Delta}\mathbf{P}$ from this equation:
$\bm{\Delta}\mathbf{P}=(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}.$ (45)
Next, we differentiate the Lagrangian function in Eq. 42 with respect to $\bm{\Lambda}$ and set it to zero:
$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Lambda}}=(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (46)
@@
$(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})\mathbf{M}\mathbf{C}_{2}=\mathbf{0}.$ (47)
@@
$\frac{1}{2}\bm{\Lambda}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}.$ (48)
@@
$(\bm{\Delta}\mathbf{P})_{\text{SPEED}}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{Q}\mathbf{M},$ (49)
@@ Appendix C Implementation Details
Table 4: Evaluation setup for multi-concept erasure. This dataset contains an erasure set with 100 celebrities and a retain set with another 100 celebrities. We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases.
@@
Figure 7: Qualitative demonstration of our erasure performance across (a) instance erasure, (b) artistic style erasure, (c) celebrity erasure, and (d) implicit concept erasure. Our method achieves precise erasure efficacy across various scenarios while exhibiting superior prior preservation. The corresponding CS is highlighted in blue, indicating that successful erasure can be achieved without pushing CS much lower, as our results demonstrate sufficient erasure at a moderate level.
@@ C.1 Experimental Setup Details
Few-concept erasure. We first compare methods on few-concept erasure, a fundamental concept erasure task, including both instance erasure and artistic style erasure following Lyu et al. (2024). For instance erasure, we prepare 80 instance templates proposed in CLIP Radford et al. (2021), such as “a photo of the {Instance}, “a drawing of the {Instance}, and “a painting of the {Instance}. For artistic style erasure, we use ChatGPT OpenAI (2022); Achiam et al. (2023) to generate 30 artistic style templates, including {Artistic} style painting of the night sky with bold strokes”, {Artistic} style landscape of rolling hills with dramatic brushwork”, and “Sunrise scene in {Artistic} style, capturing the beauty of dawn”. Following Lyu et al. (2024), we handpick the representative target and anchor concepts as the erasure set (i.e., Snoopy, Mickey, SpongeBob $\to$ ‘ ’ in instance erasure and Van Gogh, Picasso, Monet $\to$ ‘art’ in artistic style erasure) and non-target concepts for evaluation (i.e., Pikachu and Hello Kitty in instance erasure and Paul Gauguin and Caravaggio in artistic style erasure). In terms of the retain set, for instance erasure, we use a scraping script to crawl Wikipedia category pages to extract fictional character names and their page view counts with a threshold of 500,000 views from 2020.01.01 to 2023.12.31, resulting in 1,352 instances. For artistic style erasure, we use the 1,734 artistic styles collected from UCE Gandikota et al. (2024). In evaluation, we generate 10 images per template per concept, resulting in 800 and 300 images for each concept in instance erasure and artistic style erasure, respectively. Moreover, we introduce the MS-COCO captions Lin et al. (2014) to serve as general prior knowledge. In implementation, we use the first 1,000 captions to generate a total of 1000 images to compare CS and FID before and after erasure.
@@
Multi-concept erasure. We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task. Following the experiment setup from Lu et al. (2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al. (2019). This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities. The full list for both sets is presented in Table 4. We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases. In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
C.2 Erasure Configurations
Implementation of previous works. In our series of three concept erasure tasks, we mainly compare against four methods: ConAbl [^4] https://github.com/nupurkmr9/concept-ablation Kumari et al. (2023), MACE [^5] https://github.com/Shilin-LU/MACE Lu et al. (2024a), RECE [^6] https://github.com/CharlesGong12/RECE Gong et al. (2025), and UCE [^7] https://github.com/rohitgandikota/unified-concept-editing Gandikota et al. (2024), as they achieve SOTA performance across different concept erasure tasks. All the compared methods are implemented using their default configurations from the corresponding official repositories. One exception is that for MACE when erasing 50 celebrities, since it doesn’t provide an official configuration and the preserve weight varies with the number of target celebrities, we set it to $1.2\times 10^{5}$ to ensure a consistent balance between erasure and preservation.
Implementation of SPEED. In line with previous methods Kumari et al. (2023); Lu et al. (2024a); Gong et al. (2025); Gandikota et al. (2024), we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment Hertz et al. (2022). In contrast, we only edit the value matrices in the CA layers, as suggested by Wang et al. (2024b). This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images Tewel et al. (2023). In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig. 4 and 5 (further ablation comparison is provided in Appx. D.5). The augmentation times $N_{A}$ in Eq. 9 is set to 10 and the augmentation ranks $r$ in Eq. 7 is set to 1 as ablated in Appx. D.5. Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the singular vectors corresponding to the singular values below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following Fang et al. (2024). Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq. 12 to ensure invertibility.
Table 5: Complete quantitative comparison of the few-concept erasure in erasing instances from Table 1 (left). The best results are highlighted in bold, and grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
@@ C.2 Erasure Configurations
Table 6: Complete quantitative comparison of the few-concept erasure in erasing artistic styles from Table 1 (right). The best results are highlighted in bold, and grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
@@
Table 7: Quantitative comparison with SPM and SPM w/o FT in multi-concept erasure. The best results are highlighted in bold. Our method is capable of erasing up to 100 celebrities at once with low $\text{Acc}_{e}$ (%) and preserving other non-target celebrities with less appearance alteration with high $\text{Acc}_{r}$ (%), resulting in the best overall erasure performance $H_{o}$ (shaded in pink). FAIL indicates that the model collapses with noisy generations ($\text{Acc}_{e}=\text{Acc}_{r}=0.00\%$).
SPM 0.00 51.79 68.24 26.42 48.44 0.00 0.00 FAIL 26.32 52.61 0.00 0.00 FAIL 25.15 63.20
SPM w/o FT 0.00 5.08 9.68 26.38 52.23 0.00 0.00 FAIL 16.22 170.68 0.00 0.00 FAIL 14.34 245.92
@@ Appendix D Additional Experiments
D.1 More Demonstrations
We further provide qualitative visualizations of the erasure results in Fig.7, illustrating the effectiveness of our method in performing precise and targeted concept erasure across diverse scenarios. Specifically, we showcase: (a) instance erasure from Table 1 (left); (b) artistic style erasure from Table 1 (right); (c) celebrity erasure from Table 2; and (d) implicit concept erasure (e.g., nudity) from Table 8. In all cases, our method successfully removes the intended concept while preserving unrelated content, demonstrating its universal erasure applications.
We also evaluate the CLIP score (CS) before and after concept erasure to assess the erasure efficacy. As shown in Fig. 8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively). This indicates that effective erasure does not require minimizing CS to an extreme. In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure. For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure. While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge. This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
D.2 Complete Results on Few-Concept Erasure
@@
We present complete quantitative comparisons of few-concept erasure, including both CS and FID, in Table 5 and Table 6. Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
Figure 8: Comparison of CLIP scores (CS) across different erasure methods. We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue. Our method achieves successful concept erasure with moderate CS values. In contrast, RECE achieves the lowest CS by enabling more aggressive erasure. For example, removing Snoopy to the extent of producing a semantically void image, and changing Mickey into a generic person. We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.